# OmniJARVIS: Unified Vision-Language-Action Tokenization Enables Open-World Instruction Following Agents

### v1

@@ 4.2 Main Results I: Atomic Tasks
Figure 4: Evaluation results of different agents on atom tasks. The text-conditioned VPT agent (Baker et al., 2022) is from Appendix I of its paper (VPT (text)).
We select “chopping trees” [Uncaptioned image], “digging dirt” [Uncaptioned image], “mining stones” [Uncaptioned image], and “collecting wheat seeds” [Uncaptioned image] as the evaluation tasks. We directly take those short task descriptions as instructions for agents. We use text-conditioned VPT (Baker et al., 2022), Open-world Control (Cai et al., 2023a), STEVE-I (Lifshitz et al., 2023), and video-instructed GROOT (Cai et al., 2023b) as baselines. We compute the average rewards of different agents on every task in Figure 4 across 10 runs. By observing the environment and adjusting action tokens dynamically, OmniJARVIS effectively follows straightforward instructions across various scenarios. It consistently achieves a high average reward with minimal standard deviation.
4.3 Main Results II: Programmatic Tasks
To further verify the ability of OmniJARVIS to complete tasks with long sequences, we use 30 programmatic tasks to evaluate the performance of different agents. These tasks require the agent to start from an empty inventory in a new world until obtaining the final required items, which is usually a chain of atom tasks. These tasks are divided into five groups based on difficulty: wooden, food, stone, iron, and diamond. For example, the prompt for task “Obtain a diamond pickaxe” [Uncaptioned image] is “Give you nothing in the inventory, obtain a diamond pickaxe.” This task requires more game time and more complex planning for up to 10 different intermediate items (Baker et al., 2022).
Baselines are divided into two types. 1) directly outputs actions, namely the native behavior tokenizer, including STEVE-I (Lifshitz et al., 2023) and GROOT (Cai et al., 2023b). 2) using pretrained LLM as a planner to output language goals and connect the STEVE-I to execute these goals, including Zero-Shot Planner (GPT) (Huang et al., 2022a), ReAct (Yao et al., 2022), and DEPS (Wang et al., 2023c). We use success rate to evaluate the completion of tasks, that is, whether the task is completed within the specified time. The experimental results are listed in Table 1.
@@ 4.3 Main Results II: Programmatic Tasks
Table 1: Success rate of different agents on programmatic tasks. We distinguish different methods through action tokenizer, where STEVE-I (Lifshitz et al., 2023) and GROOT (Cai et al., 2023b) directly output actions, GPT (Huang et al., 2022a), ReAct (Yao et al., 2022), and DEPS (Wang et al., 2023c) output language as goals, and then use STEVE-I to output actions. OmniJARVIS uses the self-supervised discrete action tokenizer.
Method Action Tokenizer Wooden (10) Food (5) Stone (5) Iron (5) Diamond (5) Average (30)
STEVE-I Native $0.04^{\pm 0.07}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ 0.01
GROOT Native $0.05^{\pm 0.08}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ 0.02
GPT Language $0.26^{\pm 0.14}$ $0.08^{\pm 0.04}$ $0.24^{\pm 0.05}$ $0.04^{\pm 0.05}$ $0.00^{\pm 0.00}$ 0.15
ReAct Language $0.44^{\pm 0.11}$ $0.12^{\pm 0.04}$ $0.30^{\pm 0.10}$ $0.06^{\pm 0.05}$ $0.00^{\pm 0.00}$ 0.23
DEPS Language $0.78^{\pm 0.11}$ $0.12^{\pm 0.04}$ $0.68^{\pm 0.08}$ $0.16^{\pm 0.05}$ $0.04^{\pm 0.05}$ 0.43
4.4 Main Results III: Open-Ended Tasks
The open-ended tasks differ from programmatic tasks due to the lack of straightforward success criteria (Fan et al., 2022). We select the long-term creative tasks which usually need at least 5 minutes of human-playing time to finish. The task prompts can be found in Appendix D. Following image generation and video generation tasks (Heusel et al., 2017; Unterthiner et al., 2018), we take the Fréchet Sequence Distance (FSD) metrics to evaluate the correlation between agent rollout video and creative instruction. Specifically, we first ask human experts to finish the creative task prompts under randomly generated worlds and record the game-playing videos $V_{\text{human}}$. Then, we provided the task prompts for different Minecraft agents, and obtained a rollout video set $V_{\text{agent}}$. Similar to FID (Heusel et al., 2017), we used MineCLIP (Fan et al., 2022) to calculate the embedding of video clips and computed FSD for the embedding distributions of human and agent rollout videos. The analysis of the metrics can be found in Appendix B.
We further conduct creative embodied question-answering benchmarks to further evaluate the ability of the agent to complete creative instructions and grasp world knowledge. The questions answering instructions set can be found in Appendix C. The full evaluation results can be found in Table 2.
Table 2: Evaluation results on open-ended tasks, including instruction-following and embodied question-answering. The values in Instruction-Following tasks are represented by FSD scores from rollout videos, where lower scores indicate better performance. In Question-Answering tasks, the values are GPT-4 elo rating scores, with higher scores indicating better performance.
VPT STEVE-1 Voyager DEPS Ours
Instruction Following $\downarrow$ 975.93 972.79 932.16 929.52 886.25
Vicuna-7B Vicuna-13B LLaMA-2-70B ChatGPT-3.5 Ours
Question Answering $\uparrow$ 2.34 2.85 2.50 7.50 8.40
OmniJARVIS is the agent that can simultaneously complete both types of tasks and has achieved the best performance in different task sets, surpassing strong baselines including Voyager (Wang et al., 2023a) and DEPS (Wang et al., 2023c). Also, it maintains strong reasoning capability, especially on embodied question answering compared to LLM baselines (with image captions as visual context).
@@ 4.5 Insights and Analysis
Table 3: Ablation experiments on OmniJARVIS with different behavior tokenizers, vision tokenizers, and training on different interactive datasets. The first line is training on the unconditional interactive dataset, i.e., without instructions on the trajectories. OmniJARVIS with VQ-GROOT (Van Den Oord et al., 2017; Cai et al., 2023b) shows no results because of training collapse.
@@
We also evaluate training OmniJARVIS with different vision tokenization, including ImageCaptioner + LLaMA2-7B (Chen et al., 2023; Touvron et al., 2023) (basically converting the vision input into textual captions), fuyu-8b (Bavishi et al., 2023), and LLaVA-7B (Liu et al., 2024) architecture. For the ImageCaptioner+, we fix the ImageCaptioner models and only fine-tune the language model, i.e., LLaMA2-7B. We use the prediction loss of behavior tokens as the evaluation criterion, namely eval loss. We found that the model trained with LLaVA-7B architecture has the lowest evaluation loss, so we chose this model as the default model.
@@
We explore OmniJARVIS with different behavior tokenizers, including the default setting using FSQ codebook, a variant of using VQ-VAE instead of FSQ (Van Den Oord et al., 2017), and simply using sub-goal language annotation as behavior “tokens”. The evaluation results on 4 programmatic tasks are listed in Table 3. Using an FSQ tokenizer is generally better than a language goal, which confirms the advantages of using a tokenized behavior over language descriptions of behavior. The use of VQ-VAE as a quantized behavior tokenizer collapsed during the training process, so there were no results in all test tasks.
@@
We conduct an in-depth investigation of behavior tokenizers with varying codebook sizes, utilizing recommended sets of FSQ levels to approximate specified codebook dimensions (Mentzer et al., 2023) as delineated in Table 4. We evaluate performance across multiple metrics for each codebook size. Codebook Usage is quantified as the proportion of codewords utilized at least once when encoding the validation datasets. Reconstruction FSD is measured by the FSD scores derived from the MineCLIP encoder (Fan et al., 2022), processing 1,000 different demonstration videos through the FSQ-GROOT and subsequent rollout in a randomly generated environment. Additionally, we measure Resampling FSD, which is the FSD score obtained when the environment rollout is conditioned on representations sampled from the codebook. Finally, we assess the average rewards for the task “collect wood” using OmniJARVIS across varying codebook sizes. Our findings indicate that increases in codebook size correlate with enhanced average rewards and reduced FSD scores, suggesting a scalable performance in OmniJARVIS with larger codebooks.
@@
4.6 Scaling Potential of OmniJARVIS
Figure 6: Scaling potential of OmniJARVIS. Its evaluation loss continues to drop with the growth of data and model parameters. The Pearson coefficients for the 2B, 7B, and 13B models are 0.9991, 0.9999, and 0.9989 respectively.
We investigate the scaling effect (Kaplan et al., 2020; Lin et al., 2024) of data and model in OmniJARVIS by monitoring the instruction-following loss on the validation set as the amount of data increases. In addition to fine-tuning from the default LLaVA-7B, we include two additional scales: OmniJARVIS-2B (fine-tuned from LLaVA-2B with Gemma-2B language models (Hinck et al., 2024)) and OmniJARVIS-13B (fine-tuned from LLaVA-13B with LLaMA2-13B language models (Liu et al., 2024)).
The validation loss curves in Figure 6 reveal the following insights: 1) When using Omni-Tokenizer, OmniJARVIS’s instruction tuning aligns with the scaling law (Kaplan et al., 2020). All curves exhibit a log-linear decrease as the data scale increases. 2) Scaling up VLM consistently enhances performance. Notably, OmniJARVIS-7B demonstrates significantly lower losses compared to OmniJARVIS-2B. However, while improvements are consistent, the difference between OmniJARVIS-7B and OmniJARVIS-13B seems less pronounced, hinting at potential saturation when further scaling up VLM. This underscores both the scalability of OmniJARVIS and the importance of increasing data volume to match the model.
@@ 5 Related Works
Several works have explored leveraging LLMs to generate action plans for high-level tasks in embodied environments (Huang et al., 2022a; Liang et al., 2022; Brohan et al., 2022b; Zhang et al., 2023). To better perform complex planning in the environment, existing methods usually utilize chain-of-thought (Wei et al., 2022) or related methods (Yao et al., 2022). To better cope with uncertainties in open worlds, some LLM-based methods generate plans interactively with human and environmental feedback (Shinn et al., 2023; Wang et al., 2023c; Huang et al., 2022b) and retrieving from memory (Wang et al., 2023b) or internet corpus (Wang et al., 2024). However, those plans can only be executed in a language environment or require an additional controller or code executor to interact in an open world.
Figure 7: Comparative Framework of Vision-Language Action Models. (a) depicts a model where upon receiving a language instruction, actions are directly output based on the environmental state, facilitating immediate interaction with the environment at a unified frequency. Smaller models with <1B parameters like VPT (Baker et al., 2022) maintain higher frequencies (>20Hz), though their capability for complex reasoning tasks is limited. Larger models with >7B parameters such as RT-2 (Brohan et al., 2023), offer enhanced performance but operate at significantly reduced frequencies (2-3Hz). (b) illustrates a common approach utilizing large vision-language models for planning, subsequently outputting language goals (Wang et al., 2023c; Driess et al., 2023; Brohan et al., 2022a). A language-conditioned policy then translates these language goals into actions at a real-time interaction rate of 20Hz, with high-level models re-planning at less than 1Hz. This hierarchical structure balances interaction frequency and performance, while it requires language as an intermediary and additional language labels. The training process of high-level vision-language models and language-conditioned policies are separate, thus performing poorly on tasks that can not be easily connected by language. (c) (ours) mirrors the hierarchical structure of (b) but differentiates by employing a self-supervised encoder-decoder policy (Cai et al., 2023b) and FSQ quantization (Mentzer et al., 2023) as a behavior tokenizer. The upper-level vision-language models produce self-supervised behavior tokens, which are then conditioned by a policy decoder to output actions, facilitating environment interaction. The behavior tokens are injected into the training corpus of vision-language-action models, which enables end-to-end inference. This approach also eliminates the need for external language supervision and scales efficiently.
@@
In order to better utilize the knowledge inside the language model for decision-making, some methods tend to use decision datasets to fine-tune pretrained language models (Durante et al., 2024; Driess et al., 2023). Gato (Reed et al., 2022) was among the first to tokenize environment-provided actions to enable joint sequential modeling across modalities. PaLM-E (Driess et al., 2023) generates high-level instructions as texts and uses dedicated controllers to perform the task described by the output instructions. The RT series focuses more on robotics settings. Specifically, RT-1 pairs a VLM with a language-conditioned controller; RT-2 extends the VLM to directly include control tokens; RT-X generalizes to new robots and environments. A recent VLA model LEO (Huang et al., 2023) expands the perception from 2D images to 3D world and enables rich scene-level reasoning and control tasks.
@@
As LLMs have achieved remarkable reasoning results and understanding capabilities across various domains, the year 2023 has witnessed researchers adopting multiple LLM-based approaches to create open-world agents in Minecraft (Wang et al., 2023c; Zhu et al., 2023; Wang et al., 2023b, a). Some methods focus on building policies for low-level skills (Cai et al., 2023b; Lifshitz et al., 2023; Baker et al., 2022). Building upon the low-level policies to interact with the Minecraft environment, Wang et al. (2023c), Yuan et al. (2023) and Wang et al. (2023b) focus on leveraging the pre-trained language models as planners to finish programmatic tasks with in-context learning. Wang et al. (2023a) adopts the life-long learning scheme and generates code as policies to enable continual exploration. Some use expert trajectories and Minecraft corpus to fine-tune pre-trained vision language models for better embodied planning (Qin et al., 2023; Zheng et al., 2023).

### v2

@@ 4.2 Main Results I: Short-horizon Atomic Tasks
We select “chopping trees” [Uncaptioned image], “digging dirt” [Uncaptioned image], “mining stones” [Uncaptioned image], and “collecting wheat seeds” [Uncaptioned image] as the evaluation tasks. We directly take those short task descriptions as instructions for agents. We use text-conditioned VPT [2], Open-world Control [9], STEVE-I [28], and video-instructed GROOT [10] as baselines. We compute the average rewards of different agents on every task in Table 4 across 10 runs. By observing the environment and adjusting action tokens dynamically, OmniJARVIS effectively follows straightforward instructions across various scenarios. It consistently achieves a high average reward with minimal standard deviation.
Table 1: Evaluation results (rewards) on short-horizon atom tasks. The text-conditioned VPT [2] (“VPT (text)”) is from Appendix I of its paper.
Method Condition Refer to caption $\uparrow$ Refer to $\uparrow$ Refer to $\uparrow$ Refer to $\uparrow$
VPT[text] [2] Language $2.6^{\pm 0.3}$ $9.2^{\pm 0.7}$ $0.8^{\pm 0.1}$
STEVE-I [28] Language $11.0^{\pm 3.0}$ $10.0^{\pm 2.5}$ $3.2^{\pm 1.6}$ $5.1^{\pm 2.5}$
GROOT [10] Video $14.3^{\pm 4.7}$ $19.7^{\pm 8.7}$ $19.0^{\pm 11.3}$ $7.3^{\pm 0.6}$
OmniJARVIS Language $10.8^{\pm 5.2}$ $20.3^{\pm 9.2}$ $25.8^{\pm 2.9}$ $8.2^{\pm 3.6}$
Table 2: Results on open-ended instruction following.
VPT STEVE-1 Voyager DEPS Ours
Instruction Following $\downarrow$ $975.9$ $972.7$ $932.1$ $929.5$ $886.2$
Table 3: Results on open-ended question answering. We use LLM-as-judge [35] to evaluate the accuracy.
Vicuna-7B Vicuna-13B LLaMA2-70B GPT-3.5 Ours
QA $\uparrow$ $2.34$ $2.85$ $2.50$ $7.50$ $8.40$
Table 4: Success rate of different agents on long-horizon programmatic tasks.
Method Action Tokenizer Wooden (10) Food (5) Stone (5) Iron (5) Diamond (5) Average
STEVE-I [28] Native $0.04^{\pm 0.07}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.01$
GROOT [10] Native $0.05^{\pm 0.08}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.00^{\pm 0.00}$ $0.02$
GPT [23] Language $0.26^{\pm 0.14}$ $0.08^{\pm 0.04}$ $0.24^{\pm 0.05}$ $0.04^{\pm 0.05}$ $0.00^{\pm 0.00}$ $0.15$
ReAct [50] Language $0.44^{\pm 0.11}$ $0.12^{\pm 0.04}$ $0.30^{\pm 0.10}$ $0.06^{\pm 0.05}$ $0.00^{\pm 0.00}$ $0.23$
DEPS [46] Language $0.78^{\pm 0.11}$ $0.12^{\pm 0.04}$ $0.68^{\pm 0.08}$ $0.16^{\pm 0.05}$ $0.04^{\pm 0.05}$ $0.43$
4.3 Main Results II: Long-horizon Programmatic Tasks
To further verify the ability of OmniJARVIS to complete tasks with long sequences, we use 30 programmatic tasks to evaluate the performance of different agents. These tasks require the agent to start from an empty inventory in a new world until obtaining the final required items, which is usually a chain of atom tasks. These tasks are divided into five groups based on difficulty: wooden, food, stone, iron, and diamond. For example, the prompt for task “Obtain a diamond pickaxe” [Uncaptioned image] is “Give you nothing in the inventory, obtain a diamond pickaxe.” This task requires more game time and more complex planning for up to 10 different intermediate items [2]. We list all programmatic tasks and its corresponding instructions in the Appendix C.1.
Baselines are divided into two types: 1) directly outputs actions, namely the native behavior tokenizer, including STEVE-I [28] and GROOT [10]. 2) using pretrained LLM as a planner to output language goals and connect the STEVE-I to execute these goals, including Zero-Shot Planner (GPT) [23], ReAct [50], and DEPS [46]. We use success rate to evaluate the completion of tasks, that is, whether the task is completed within the specified time. The experimental results are listed in Table 6.
@@ 4.3 Main Results II: Long-horizon Programmatic Tasks
4.4 Main Results III: Open-ended Question-Answering and Instruction Following Tasks
The open-ended tasks differ from programmatic tasks due to the lack of straightforward success criteria [16]. We select the long-term open-ended tasks which usually need at least 5 minutes of human-playing time to finish. The task prompts can be found in Appendix C.3. Following image generation and video generation tasks [19, 41], we take the Fréchet Sequence Distance (FSD) metrics to evaluate the correlation between agent rollout video and creative instruction. Specifically, we first ask human experts to finish the creative task prompts under randomly generated worlds and record the game-playing videos $V_{\text{human}}$. Then, we provided the task prompts for different Minecraft agents, and obtained a rollout video set $V_{\text{agent}}$. Similar to FID [19], we used MineCLIP [16] to calculate the embedding of video clips and computed FSD for the embedding distributions of human and agent rollout videos. The analysis of the metrics can be found in Appendix B.
We further conduct open-ended embodied question-answering benchmarks to evaluate the ability of the agent to complete open-ended instructions and grasp world knowledge. The questions answering instructions set can be found in Appendix C.2. The evaluation results can be found in Table 4 and Table 4. OmniJARVIS is the agent that can simultaneously complete both types of tasks and has achieved the best performance in different task sets, surpassing strong baselines including Voyager [44] and DEPS [46]. Also, it maintains strong reasoning capability, especially on embodied question answering compared to LLM baselines (with image captions as visual context).
@@ 4.5 Insights and Analysis
Figure 4: Ablation experiments on OmniJARVIS with different behavior tokenizers, vision tokenizers, and training on different interactive datasets. The first line is training on the unconditional interactive dataset, i.e., without instructions on the trajectories. OmniJARVIS with VQ-GROOT [42, 10] shows no results because of training collapse.
@@
Figure 5: Scaling potential of OmniJARVIS. Its evaluation loss continues to drop with the growth of data and model parameters. The Pearson coefficients for the 2B, 7B, and 13B models are 0.9991, 0.9999, and 0.9989.
@@
We also evaluate training OmniJARVIS with different vision tokenization, including ImageCaptioner + LLaMA2-7B [12, 40] (basically converting the vision input into textual captions), fuyu-8b [3], and LLaVA-7B [31] architecture. For the ImageCaptioner+, we fix the ImageCaptioner models and only fine-tune the language model, i.e., LLaMA2-7B. We use the prediction loss of behavior tokens as the evaluation criterion, namely eval loss. We found that the model trained with LLaVA-7B architecture has the lowest evaluation loss, so we chose this model as the default model.
@@
We explore OmniJARVIS with different behavior tokenizers, including the default setting using FSQ codebook, a variant of using VQ-VAE instead of FSQ [42], and simply using sub-goal language annotation as behavior “tokens”. The evaluation results on 4 programmatic tasks are listed in Table 5. Using an FSQ tokenizer is generally better than a language goal, which confirms the advantages of using a tokenized behavior over language descriptions of behavior. The use of VQ-VAE as a quantized behavior tokenizer collapsed during the training process, so there were no results in all test tasks.
@@
We conduct an in-depth investigation of behavior tokenizers with varying codebook sizes, utilizing recommended sets of FSQ levels to approximate specified codebook dimensions [34] as delineated in Table 5. We evaluate performance across multiple metrics for each codebook size. Codebook Usage is quantified as the proportion of codewords utilized at least once when encoding the validation datasets. Reconstruction FSD is measured by the FSD scores derived from the MineCLIP encoder [16], processing 1,000 different demonstration videos through the FSQ-GROOT and subsequent rollout in a randomly generated environment. Additionally, we measure Resampling FSD, which is the FSD score obtained when the environment rollout is conditioned on representations sampled from the codebook. Finally, we assess the average rewards for the task “collect wood” using OmniJARVIS across varying codebook sizes. Our findings indicate that increases in codebook size correlate with enhanced average rewards and reduced FSD scores, suggesting a scalable performance in OmniJARVIS with larger codebooks.
@@
Figure 7: OmniJARVIS plays Montezuma’s Revenge and gets a reward of 3600.
4.6 Generalization and Scaling Potential of OmniJARVIS
We first explore adapting OmniJARVIS to the Atari game Montezuma’s Revenge. We created a dataset from 500 episodes played by an agent trained with Random Network Distillation [8], supplemented by random actions in early frames to enhance diversity. This dataset contains 1,823,699 transitions. We then trained the FSQ-GROOT tokenizer on this new dataset and subsequently trained OmniJARVIS on the tokenized data. The finetuned OmniJARVIS achieved a score of 3600 in Montezuma’s Revenge, indicating promising transferability. A rollout trajectory is in Figure 7.
We also investigate the scaling effect [25, 30] of data and model in OmniJARVIS by monitoring the instruction-following loss on the validation set as the amount of data increases. In addition to fine-tuning from the default LLaVA-7B, we include two additional scales: OmniJARVIS-2B (fine-tuned from LLaVA-2B with Gemma-2B language models [20]) and OmniJARVIS-13B (fine-tuned from LLaVA-13B with LLaMA2-13B language models [31]).
The validation loss curves in Figure 5 reveal the following insights: 1) When using Omni-Tokenizer, OmniJARVIS’s instruction tuning aligns with the scaling law [25]. All curves exhibit a log-linear decrease as the data scale increases. 2) Scaling up VLM consistently enhances performance. Notably, OmniJARVIS-7B demonstrates significantly lower losses compared to OmniJARVIS-2B. However, while improvements are consistent, the difference between OmniJARVIS-7B and OmniJARVIS-13B seems less pronounced, hinting at potential saturation when further scaling up VLM. This underscores both the scalability of OmniJARVIS and the importance of increasing data volume to match the model.
@@ 5 Related Works
Several works have explored leveraging LLMs to generate action plans for high-level tasks in embodied environments [23, 27, 5, 52]. To better perform complex planning in the environment, existing methods usually utilize chain-of-thought [49] or related methods [50]. To better cope with uncertainties in open worlds, some LLM-based methods generate plans interactively with human and environmental feedback [39, 46, 24] and retrieving from memory [47] or internet corpus [48]. However, those plans can only be executed in a language environment or require an additional controller or code executor to interact in an open world.
Figure 8: Comparative Framework of Vision-Language Action Models. (a) depicts a model where upon receiving a language instruction, actions are directly output based on the environmental state, facilitating immediate interaction with the environment at a unified frequency. Smaller models with <1B parameters like VPT [2] maintain higher frequencies (>20Hz), though their capability for complex reasoning tasks is limited. Larger models with >7B parameters such as RT-2 [6], offer enhanced performance but operate at significantly reduced frequencies (2-3Hz). (b) illustrates a common approach utilizing large vision-language models for planning, subsequently outputting language goals [46, 14, 4]. A language-conditioned policy then translates these language goals into actions at a real-time interaction rate of 20Hz, with high-level models re-planning at less than 1Hz. This hierarchical structure balances interaction frequency and performance, while it requires language as an intermediary and additional language labels. The training process of high-level vision-language models and language-conditioned policies are separate, thus performing poorly on tasks that can not be easily connected by language. (c) (ours) mirrors the hierarchical structure of (b) but differentiates by employing a self-supervised encoder-decoder policy [10] and FSQ quantization [34] as a behavior tokenizer. The upper-level vision-language models produce self-supervised behavior tokens, which are then conditioned by a policy decoder to output actions, facilitating environment interaction. The behavior tokens are injected into the training corpus of vision-language-action models, which enables end-to-end inference. This approach also eliminates the need for external language supervision and scales efficiently.
@@
In order to better utilize the knowledge inside the language model for decision-making, some methods tend to use decision datasets to fine-tune pretrained language models [15, 14]. Gato [38] was among the first to tokenize environment-provided actions to enable joint sequential modeling across modalities. PaLM-E [14] generates high-level instructions as texts and uses dedicated controllers to perform the task described by the output instructions. The RT series focuses more on robotics settings. Specifically, RT-1 pairs a VLM with a language-conditioned controller; RT-2 extends the VLM to directly include control tokens; RT-X generalizes to new robots and environments. A recent VLA model LEO [22] expands the perception from 2D images to 3D world and enables rich scene-level reasoning and control tasks.
@@
As LLMs have achieved remarkable reasoning results and understanding capabilities across various domains, the year 2023 has witnessed researchers adopting multiple LLM-based approaches to create open-world agents in Minecraft [46, 55, 47, 44]. Some methods focus on building policies for low-level skills [10, 28, 2]. Building upon the low-level policies to interact with the Minecraft environment, Wang et al. [46], Yuan et al. [51] and Wang et al. [47] focus on leveraging the pre-trained language models as planners to finish programmatic tasks with in-context learning. Wang et al. [44] adopts the life-long learning scheme and generates code as policies to enable continual exploration. Some use expert trajectories and Minecraft corpus to fine-tune pre-trained vision language models for better embodied planning [36, 54].